a. Using the data below as of 12/31/17 and the format found in Exhibit 4.1, create a Balance Sheet for XYZ Hospital for 12/31/17 Based upon your Balance Sheet, answer the following in the space provided:
Accounts payable 401,760
Accrued expenses 485,520
Cash 968,160
Inventory 295,600
Long-term debt 6,825,760
Long-term investments 3,844,720
Net patient accounts receivable 2,280,720 b. How much does the hospital own?
Net property and equipment 4,196,000 c. What % of the hospital's assets are current?
Notes payable 346,720 d. How much does the hospital owe to others?
Other long-term liabilities 420,000 e. What are current assets minus current liabilities?
Short-term investments 800,000 f. What is the hospital's net assets (equity)?
g. If the hospital earned $250,000 in 2017, what was the net assets (equity) balance on 12/31/16?
h. If the net assets (equity) balance was $375,000 on 12/31/16, what was the hospital's net income in 2017?
** use the space below to create the Balance Sheet:
Answer #1

XYZ Hospital

Balance sheet

December 31, 2017

Assets

Current assets:

Cash

968,160

Net patient accounts receivable

2,280,720

Short-term investments

800,000

Inventory

295,600

Total current assets

4344480

Non-current assets:

Long-term investments

3,844,720

Net property and equipment

4,196,000

Total noncurrent assets

8040720

Total assets

12385200

Liabilities and equity

Current liabilities

Accounts payable

401,760

Accrued expenses

485,520

Total current liabilities

887280

Long-term liabilities

Long-term debt

6,825,760

Other long-term liabilities

420,000

Total long term liabilities

7245760

Total liabilities

8133040

Equity

4252160

Total liabilities and equity

12385200

Part B

Hospital own $12385200

This will include the hospital’s total assets.

Part C

The hospital's 35.08% or 35% (4344480/12385200) of assets are current.

% of current assets = current assets / total assets *100

Part D

The hospital $8133040 owe to others.

This will include the hospital’s total liabilities

Part E

Current assets minus current liabilities is referred as the working capital.

Current assets minus current liabilities = 4344480-887280 = $3457200

Part F

Hospital's net assets (equity) = Total assets – total liabilities = 12385200-8133040 = $4252160

Part G

the net assets (equity) balance on 12/31/16 = the net assets (equity) balance on 12/31/17 – net income in 2017 = 4252160-250000 = 4002160

Part H

Net income in 2017 = net assets (equity) balance on 12/31/17 - net assets (equity) balance on 12/31/16 = 4252160-375000 = $3877160